OpenStudy (anonymous):

3^2x = 27 solve for x son!

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=3%5E2x%3D27

OpenStudy (anonymous):

Scroll down, and you have the method to solve it :D

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

you aint playing are ya?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

no u tell me the answer

OpenStudy (anonymous):

Don't worry, it is wolfram alpha, you can go on the website yourself and type your problem in

OpenStudy (mertsj):

Please note that 27 is a power of 3 so write: \[3^{2x}=3^3\]

OpenStudy (anonymous):

really?

OpenStudy (mertsj):

And so we know that 2x =3

OpenStudy (anonymous):

I'll send you a picture of it @tomtom777

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I suggest you use the website @tomtom777

OpenStudy (mertsj):

Is it 9x or 3^(2x) ????

OpenStudy (anonymous):

It's 3^2x

OpenStudy (mertsj):

so 9x ?????

OpenStudy (nathan917):

(x-2)/3 = 27 / (x-2) (<-- better way of writing this) cross-multiply!! (x-2)*(x-2) = 27 * 3 you could multiply out both sides and solve the quadratic equation but this one seems clear without it... (x-2) squared = 9 squared !!! so (x-2) = + 9 or -9 so x = 11 or -7 how about this?

OpenStudy (anonymous):

@nathan917 I think you made a mistake

OpenStudy (nathan917):

what would that be?

OpenStudy (anonymous):

..... what im so confused now

OpenStudy (anonymous):

The question @tomtom777 wrote was 3^2x=27. So,

OpenStudy (mertsj):

@tomtom777 What is 3^2 ???

OpenStudy (nathan917):

k

OpenStudy (mertsj):